OpenStudy (anonymous):

What are the roots of the equation, -3x^2 +8x+10=0

hartnn (hartnn):

do you know Quadratic formula ??

OpenStudy (anonymous):

I can't remember it

hartnn (hartnn):

Compare your quadratic equation with \(ax^2+bx+c=0\) find a,b,c then the two roots of x are: \(\huge{x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}}\)

OpenStudy (anonymous):

Thanks!

hartnn (hartnn):

welcome ^_^ tell us what 2 values you get, so that we can verify :)